OpenStudy (anonymous):

Solve using Cramer's Rule. 2x + 2y = –8 7x + 4y = –1

OpenStudy (anonymous):

4x+4y=-16 - 7x+4y= -1 ------------- -3x = -15 x = 5 2(5) +2y=-8 10+2y=-8 2y=-18 y=-9

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

can you help with another?

OpenStudy (anonymous):

Sure. Which is it?

OpenStudy (anonymous):

Solve the system of equations. –5x + 3y + 2z = 19 x + 4y + 3z = –27 15x – 9y – 6z = –55

OpenStudy (anonymous):

Hold on. Let me see.

OpenStudy (anonymous):

I'm sorry... I can't help much there! But I'm sure someone will be able if you post this question up

OpenStudy (anonymous):

ok thanks

OpenStudy (mathmate):

Did you want it solved by Cramer's rule?

OpenStudy (mathmate):

By Cramer's rule it would be: 2x + 2y = –8 7x + 4y = –1 \[D=\det \left[\begin{matrix}2 & 2 \\ 7 & 4\end{matrix}\right] = 2*4-7*2=-6\] \[Dx=\det \left[\begin{matrix}-8 & 2 \\ -1 & 4\end{matrix}\right] = -32+2 = -30\] and \[Dy=\det \left[\begin{matrix}2 & -8 \\ 7 & -1\end{matrix}\right] = -2+56 = 54\] So x=Dx/D=-30/(-6)=5 y=Dy/D=54/(-6)=-9